OpenStudy (anonymous):

i'll fan & MEDAL!!!! If p − 1 is a factor of p^4 + p^2 + p − k, the value of k is _________.

OpenStudy (amoodarya):

\[p^4 + p^2 + p − k=(p-1)Q(p)\\\] if you put p=1 into rhs,lhs so you will find k

OpenStudy (anonymous):

what is rhs & lhs

OpenStudy (amoodarya):

left hand side =right hand side l h s = r h s

OpenStudy (anonymous):

(p−1)Q(p) im confused with this equation. Im suppose to use synthetic division

OpenStudy (anonymous):

Do you know how to use synthetic division? It's sort of a shortcut.

OpenStudy (amoodarya):

it is easy it mean p^4 + p^2 + p − k dive to (p-1) it means too , no remainder

OpenStudy (anonymous):

yes but for this equation i dont get what the value of k could be

OpenStudy (anonymous):

so the value is 1

OpenStudy (anonymous):

Wait how can you what? Is this an equation?

OpenStudy (amoodarya):

put p=1 \[p^4 + p^2 + p − k=(p-1)Q(p)\\ 1+1+1-k=0 Q(1)=0\\so k= ?\]

OpenStudy (anonymous):

Ohm

OpenStudy (anonymous):

is it 4

OpenStudy (amoodarya):

1+1+1-k=0 k=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Thank you so much